Suppose the marginal propensity to consume is 0.8 and the tax rate is 0.25 and all other components of aggregate expenditures are determined outside the model. If the president wants to increase income by 500, her advisers would suggest that she increases government spending by: A. 50. O B. 100 C. 200 D. 250
Answer #1

B. 100

Explanation: Spending multiplier = 1/(1-MPC) = 1/(1-0.8) = 1/0.2 = 5

This means an increase in government expenses by $1 would increase income by $5.

So, in order to increase income by $500, the government spending needs to increase by $500/5 = $100.